OpenStudy (anonymous):

A buffer is created by combining 140.0 ml of 0.25M HCHO2 with 75.0 ml of 0.20M NaOH .Determine the of the buffer.

OpenStudy (anonymous):

*determine the PH

OpenStudy (anonymous):

ok then..?

OpenStudy (anonymous):

u knw how to calculate pH by concentration?

OpenStudy (anonymous):

is that PH =- log [H30+]..?

OpenStudy (anonymous):

nah i m ot able to take log (24x 10^-1) i get this by N1V1 +N2V2 = NV N normality V volume 140x0.25 + 0.20x75=N215 35+15=215V 50/215=N 0.23 ssmthing is coming

OpenStudy (anonymous):

i m gettin pH=1.38

OpenStudy (anonymous):

a minute

OpenStudy (anonymous):

it's wrong..=(

OpenStudy (anonymous):

a by rounding i take N=24 x 10^-1 pH=log(24 x 10^-1) =log(6x4x10^-1) =log6 + log 4 + log 10^-1 =0.78+0.60 + 1-log 10 =1.38 + 0

OpenStudy (anonymous):

hmm.. i got this but (:

OpenStudy (anonymous):

are u really sure..? sorry but this is my last attempt..on9 quiz..=(

OpenStudy (anonymous):

saorry i m gettin last try ans is 1.0458 this is the final ans

OpenStudy (anonymous):

can u tell me the options?

OpenStudy (anonymous):

its no option..=(

OpenStudy (anonymous):

1.05 final try this one

OpenStudy (anonymous):

can u give me the link og quiz if u dont mind

OpenStudy (anonymous):

in need to login..

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

nah i was jut aking gime me the site name if u remeber

OpenStudy (anonymous):

mylabsplus..=)

OpenStudy (anonymous):

heena..if u dont mind, how to clculate pH of N2H4 when concentration is 0.25M?

OpenStudy (anonymous):

simple N2H4 there the charge is 1 so to caluclate normality first so multiply concentration given in molarity with charge now pH=log concentration pH=log (25 x 10^-2) pH=2 - log 25 pH=2-log(5x5) pH=2-log5 +log 5 = 2-0.7+0.7 =2-1.4 =0.6

OpenStudy (anonymous):

alway remeber before taking normality always mulitply molarity x charge like in H2SO4 charge is 2 so now concentration change ok

OpenStudy (anonymous):

normality = pH?

OpenStudy (anonymous):

how to know the charge??

OpenStudy (anonymous):

sorry but it says that N2H4 is a weak base..

OpenStudy (anonymous):

pOH=log concentration pOH=log (25 x 10^-2) pOH=2 - log 25 pOH=2-log(5x5) pOH=2-log5 +log 5 = 2-0.7+0.7 =2-1.4 pOH=0.6 now we know pH-pOH=14

OpenStudy (anonymous):

sry wat hapening me i dunno u r right N2H4 is base so we will get pOH now by completing above que u ll get pH

OpenStudy (jfraser):

You start with a "large" amount of weak acid (formic acid, HCHO2), and add to it a small amount of strong base (NaOH) which will neutralize SOME of the weak acid into the formate ion (CHO2^-1). You have to find the moles of base added, which will tell you the moles of formate ion formed. From that, find the moles of acid that were used, and the moles of acid that are left over. Turn those moles into concentrations, and use the equation that determines the pH of a buffer: \[pH = pK_{A} * \log(\frac{[Base]}{[Acid]})\]

OpenStudy (anonymous):

but here u are not given pka

OpenStudy (jfraser):

look up the KA of formic acid

OpenStudy (anonymous):

i think i m lacking smwhere if u dont mind can u post the formula how to calculate KA

OpenStudy (anonymous):

The pKa of formic acid is 1.8 x 10–4. Now apply Henderson's equation, pH=pKa+log(salt/acid) salt=meq of base acid=meq of acid-meq of base.